# subfigure0(#1)

### v2

@@ Abstract.
Key words and phrases:
ball graph, disk graph, vertex placement, reliability, area coverage, autonomous swarm, satellite swarm, swarm robotics, formation planning
1991 Mathematics Subject Classification:
05C85 (primary); 68R10, 93B51 (secondary)
Implementation in Python 3 of the algorithms described herein is publicly available on GitHub at https://github.com/calum-buchanan/UDG-Neighborhoods-Reliability-AreaCoverage.
1. Introduction
@@
One method to keep agents spread out is to ensure some minimum distance between any given pair. The main contribution of this paper is an efficient algorithm to enumerate all possible neighborhoods that a vertex $v$ could have when added to a unit disk graph, under the constraint that no other vertices are within some fixed Euclidean distance $b$ of $v$ (Algorithm 1 of Section 3.1). We use this algorithm to move vertices one at a time, under the same constraint, to locations which maximize the reliability of the resulting graph. An alternative method is to use a spring layout- (or Fruchterman-Reingold)-type algorithm, treating vertices like similarly charged particles which repel and edges like springs which attract or repel based on the resting length; we provide a modified spring layout algorithm which avoids destroying edges in a unit ball graph in Section 3.2.
The remainder of this paper is organized as follows. First, in Section 2, we review an algorithm from our preliminary conference paper [8] to enumerate all possible neighborhoods for a vertex added to a unit ball graph. This can be used to move a vertex in a unit ball graph to a location which maximizes the network reliability of the resulting graph. In Section 3, we consider both reliability and area coverage. In Section 3.1, we describe our main algorithm in terms of unit disk graphs, noting that this algorithm can be generalized to unit ball graphs as well. In Section 3.2, we provide a variation of a classical algorithm of Fruchterman and Reingold [19] to space out vertices in a ball graph without destroying edges. In Section 4, we compare the effectiveness of the algorithms in Sections 3.1 and 3.2 in producing unit disk graphs with high reliability and evenly spread vertices over a region. The comparison shows the former algorithm to be particularly effective. We conclude with some open questions and future directions in Section 5.
1.1. Definitions and notations
@@ 1.2. Problem statement
2. Radial sweeps and ball graphs
We begin with preliminaries on radial sweep algorithms as they apply to disk and ball graphs. We outline an algorithm from our preliminary paper [8], using radial sweeps to list all possible neighborhoods that a new vertex could have when added to a disk or ball graph. This will provide a basis for our main algorithm, described in Section 3.
2.1. Radial Sweeps
@@
$\theta_{\mathrm{min}}$$\theta_{\mathrm{max}}$$v$$w$
Figure 1. A circle of radius $1$ revolves in a counterclockwise motion around a fixed point $v$ on its boundary in a radial sweep. A point $w$ enters and exits the circle at the angles $\theta_{\mathrm{min}}$ and $\theta_{\mathrm{max}}$, respectively. (Adapted from [8].)
Performing these trigonometric calculations for each vertex $v$ in $V$ and each vertex $w$ in $V-v$ with $\|v-w\|_{2}<2$, one can determine all subsets $N$ of $V$ for which some radius-$1$ circle contains all of the points in $N$ and none in $V-N$ as follows. Each time a point $w\in V$ enters or exits the circle sweeping around a point $v\in V$, we record the subset $S$ of $V$ contained in the interior this circle with $v$ and $w$ on its boundary. By shifting and/or rotating the circle slightly, one can obtain a unit circle with any of $S$, $S\cup v$, $S\cup w$, or $S\cup\{v,w\}$ in its interior. All of these are subsets $N$ of $V$ as described. On the other hand, if $N$ is the subset of $V$ contained in a given radius-$1$ circle, then the circle can be shifted until two points in $V$ lie on its boundary. Thus, after performing a radial sweep around each vertex in $V$, we obtain in this manner all such subsets $N$ of $V$.
@@
2.2. Unit disk and unit ball graphs
@@
Given the dynamic nature of swarm formations, it is natural to ask how the communication network changes with the formation. Generally, one might like to know which changes to the graph structure of a unit ball graph $G$ are obtainable by repositioning vertices. This problem is difficult in general, as the recognition problem for unit ball graphs is NP-hard [7]. However, using the algorithm described in Section 2.1, we are able to enumerate in polynomial time all of the unit ball graphs obtainable by moving a single vertex in $G$. This can be modeled by deleting a vertex $v$ from $G$ and adding it back in a new location. The algorithm summarized in Section 2.1 can be used to list all of the possible neighborhoods that $v$ could have.
3. Reliability and area coverage
@@
In the context of many a swarm mission, like satellite imaging of a region, such a formation may not be desirable. A natural question then arises: which unit disk graphs have evenly distributed vertices over a given area and have high reliability? In the spirit of our first algorithm, we propose to add or move vertices one at a time, maximizing reliability at each iteration, but imposing the constraint that no vertex be within some fixed distance $b$ of the vertex being added or moved. We present such an algorithm in Section 3.1, and we propose an alternative solution to the area coverage problem using a modified spring layout algorithm in Section 3.2.
3.1. Finding neighborhoods for a vertex with a buffer
@@
There are a number of ways that one might handle this problem. We propose a method whereby a buffer distance is imposed around each vertex as it is added to the unit disk graph $G$. For a value $b$ with $10<b<$ 1, we note that a vertex which has no neighbors within distance $b$ must have all of its neighbors contained in the $(b,1)$-annulus centered at that vertex (see Figure 2).
@@
Figure 2. The neighbors of a vertex $v$ added with a buffer to a disk graph are contained within an annulus with no vertices in its inner circle. The buffer is depicted by a shaded region, and existing edges in the graph before the addition of $v$ are depicted by dashed edges.
In other words, the possible neighborhoods for a new vertex added with a buffer are precisely the subsets $S$ of the vertex set $V$ of $G$ such that $S$ is contained in a $(b,1)$-annulus and all points in $V-S$ lie outside of the radius-$1$ circle. We use a modified radial sweep algorithm to determine all such subsets $S$ of $V$. This subsection is devoted to proving the following.
@@
Let $G$ be a unit disk graph with vertex set $V$, and let $10<b<$ 1. The list of possible neighborhoods for a vertex $v$ added to $G$ such that $\|v-w\|_{2}>b$ for all $w\in V$ can be obtained in $\mathcal{O}(n\Delta)$ time, where $\Delta$ is the maximum cardinality of a subset of $V$ contained in a radius-$2$ circle.
@@
Algorithm 1 Enumerate neighborhoods for a new vertex added to a unit disk graph with no other vertices within distance $b$ ($10<b<$ 1).
@@
7:         $L_{1}$leftarrow $\textsc{OuterSweep}(V,v,b)$
@@
11:         $L_{2}$leftarrow $\textsc{InnerSweep}(V,v,b)$
@@
[…]
6:     for $w\in V-v$ with $\|v-w\|_{2}<2$ do $\triangleright$ $w$ will enter & exit the sweep
7:         if $1-b<\|v-w\|_{2}<1+b$ then $\triangleright$ $w$ will enter & exit both circles
8:              $\theta_{\mathrm{min}}$leftarrow angle $w$ enters outer circle
9:              $\theta_{\mathrm{dis}}$leftarrow angle $w$ enters inner circle
10:              $\theta_{\mathrm{re}}$leftarrow angle $w$ exits inner circle
11:              $\theta_{\mathrm{max}}$leftarrow angle $w$ exits outer circle
12:         $\triangleright$ $w$ will only enter & exit the outer circle
13:              $\theta_{\mathrm{min}}$leftarrow angle $w$ enters outer circle
14:              $\theta_{\mathrm{dis}}$, $\theta_{\mathrm{re}}$leftarrow None
15:              $\theta_{\mathrm{max}}$leftarrow angle $w$ exits outer circle
16:         end if
17:         add $(w,\theta_{\mathrm{min}},\theta_{\mathrm{dis}},\theta_{\mathrm{re}},\theta_{\mathrm{max}})$ to $L$
18:     end for
19:     return $L$
20:end function
@@
[…]
6:     for $w\in V-p$ with $\|v-w\|_{2}<1+b$ do $\triangleright$ $w$ will enter & exit the sweep
7:         if $1-b<\|v-w\|_{2}<2b$ then $\triangleright$ $w$ will enter & exit both circles
8:              $\theta_{\mathrm{min}}$leftarrow angle $w$ enters outer circle
9:              $\theta_{\mathrm{dis}}$leftarrow angle $w$ enters inner circle
10:              $\theta_{\mathrm{re}}$leftarrow angle $w$ exits inner circle
11:              $\theta_{\mathrm{max}}$leftarrow angle $w$ exits outer circle
12:         else if $\|v-w\|_{2}>2b$ then $\triangleright$ $w$ will only enter & exit outer circle
13:              $\theta_{\mathrm{min}}$leftarrow angle $w$ enters outer circle
14:              $\theta_{\mathrm{dis}}$, $\theta_{\mathrm{re}}$leftarrow None
15:              $\theta_{\mathrm{max}}$leftarrow angle $w$ exits outer circle
16:         else if $\|v-w\|_{2}<1-b$ then $\triangleright$ $w$ will only enter & exit inner circle
17:              $\theta_{\mathrm{min}}$, $\theta_{\mathrm{max}}$leftarrow None
18:              $\theta_{\mathrm{dis}}$leftarrow angle $w$ enters inner circle
19:              $\theta_{\mathrm{re}}$leftarrow angle $w$ exits inner circle
20:         end if
21:         add $(w,\theta_{\mathrm{min}},\theta_{\mathrm{dis}},\theta_{\mathrm{re}},\theta_{\mathrm{max}})$ to $L$
22:     end for
23:     return $L$
24:end function
First, we sweep a $(b,1)$-annulus around a fixed point $v\in V$ on its outer boundary (see Algorithm 2). Consider the $(b,1)$-annulus whose center is at $v+\left[\begin{matrix}1&0\end{matrix}\right]^{\mathsf{T}}$. We imagine a full counterclockwise rotation of the annulus around the fixed point $v$. Points $w\in V$ with $1-b<\|v-w\|_{2}<1+b$ will enter and exit both the inner and outer circles of the annulus as it rotates. Using trigonometric equations similar to those described in Section 2.1, we record the angles of rotation for each of these four moments: let $\theta_{\mathrm{min}}$ denote the angle at which $w$ enters the outer circle, $\theta_{\mathrm{dis}}$ the angle at which $w$ enters the inner circle (and disappears from the annulus), $\theta_{\mathrm{re}}$ the angle at which $w$ exits the inner circle (and reappears in the annulus), and $\theta_{\mathrm{max}}$ the angle at which $w$ exits the outer circle. We add the $5$-tuple $(w,\theta_{\mathrm{min}},\theta_{\mathrm{dis}},\theta_{\mathrm{re}},\theta_{\mathrm{max}})$ to a list $L$. Points $w\in V$ with $\|v-w\|_{2}<1-b$ or $\|v-w\|_{2}>1+b$ will enter and exit only the outer circle of the annulus as it rotates about $v$. In this case, we set $\theta_{\mathrm{dis}}=\mathrm{None}$ and $\theta_{\mathrm{re}}=\mathrm{None}$.
@@
5:     $M$leftarrow [][\ ][ ]
6:     $R$leftarrow $\textsc{InitialSet}(L)$triangleright See Algorithm 5 in Appendix B
[…]
8:     for $(w,\theta_{\mathrm{min}},\theta_{\mathrm{dis}},\theta_{\mathrm{re}},\theta_{\mathrm{max}})\in L$ and  do
9:         for $\theta\in\{\theta_{\mathrm{min}},\theta_{\mathrm{dis}},\theta_{\mathrm{re}},\theta_{\mathrm{max}}\}$ do
10:              if $\theta$ is not None then
11:                  add $(w,\theta,\mathrm{type})$ to $M$, where $\mathrm{type}\in\{\min,\mathrm{dis},\mathrm{re},\max\}$
12:              end if
13:         end for
14:         if $\theta_{\mathrm{re}}<\theta_{\mathrm{dis}}$ then $\triangleright$ $w$ is bad when sweep begins
15:              add $(w,\theta_{\mathrm{dis}},\mathrm{type}=\mathrm{dis})$ to $B$
16:         end if
17:     end for
18:     Sort $M$ by angles $\theta$ in increasing order
[…]
20:     for $(w,\theta,\mathrm{type})\in M$ do
21:         if $\mathrm{type}=\max$ then remove $w$ from $R$
22:              if $B=\emptyset$ then add $R$ and $R\cup w$ to $\mathrm{Sets}$
23:              end if
24:         else if $\mathrm{type}=\min$ then add $w$ to $R$
25:              if $B=\emptyset$ then add $R$ and $R-w$ to $\mathrm{Sets}$
@@
27:         else if $\mathrm{type}=\mathrm{dis}$ then remove $w$ from $R$
28:              if $B=\emptyset$ then add $R\cup w$ to $\mathrm{Sets}$
@@
30:              add $w$ to $B$
31:         else if $\mathrm{type}=\mathrm{re}$ then add $w$ to $R$
32:              remove $w$ from $B$
33:              if $B=\emptyset$ then add $R$ to $\mathrm{Sets}$
34:              end if
35:         end if
36:     end for
37:     return $\mathrm{Sets}$
38:end function
We now describe how to obtain the sets $S$ of vertices contained in a $(b,1)$-annulus with $V-S$ outside of its outer boundary (see Algorithm 4). By Proposition 2, it suffices to find the sets $S$ contained in a $(b,1)$-annulus with two points on its boundary (keeping track of whether they are on the outer or inner boundary) and all other points in $V-S$ outside of the outer boundary. One of the points on the boundary will be the point $v$ we are sweeping around, and the other will be a point that is entering or exiting the sweep.
@@
From $L_{v}$, we make a new list $M_{v}$ of tuples $(w,\theta,\mathrm{type})$ to record the angles $\theta$, in increasing order, at which points $w$ enter and exit the annulus during the sweep, along with whether it is entering or exiting and which boundary it is passing through. Specifically, $\mathrm{type}$ is one of ‘$\min$’, ‘$\mathrm{dis}$’, ‘$\mathrm{re}$’, or ‘$\max$’ depending on whether $\theta$ is the angle at which $w$ enters the outer circle, enters the inner circle, exits the inner circle, or exits the outer circle of the annulus, respectively.
Now, we go through the tuples in $M_{v}$ in increasing order of $\theta$. If a vertex $w$ exits the outer circle (that is, if the tuple $(w,\theta,\mathrm{type})$ has type ‘$\max$’), we remove $w$ from $R$. We then check if the set of bad vertices $B$ is empty and, if it is, add both $R$ and $R\cup w$ to a list $\mathrm{Sets}$ which will be the output of Algorithm 4. If a vertex $w$ enters the outer circle (if the tuple has type ‘$\min$’), we add it to $R$, and if $B=\emptyset$, we add $R$ and $R-w$ to $\mathrm{Sets}$. If a vertex $w$ enters the inner circle (if the tuple has type ‘$\mathrm{dis}$’), we remove $w$ from $R$, check if $B=\emptyset$, and, if it is, add $R\cup w$ to $\mathrm{Sets}$. Then, we add $w$ to $B$. Finally, if a vertex $w$ exits the inner circle (if the tuple has type ‘$\mathrm{re}$’), we add it to $R$ and remove it from $B$. If now $B=\emptyset$, we add $R$ to $\mathrm{Sets}$.
@@
Performing Algorithms 2 and 3 each $n=|V|$ times, by Proposition 2 we obtain all subsets $S$ of $V$ for which there exists a $(b,1)$-annulus containing $S$ and with every point in $V-S$ outside of the radius-$1$ circle in $\mathcal{O}(n^{2})$ time. This is summarized in Algorithm 1.
@@
By adjusting the value of $b$, we can use this algorithm to fill in a region with a reliable graph whose vertices are evenly spaced apart. In Section 4, we will compare this algorithm to another method for spreading out vertices without decreasing reliability which we presently describe.
3.2. Fruchterman-Reingold for ball graphs
We now describe an alternative approach for modifying a unit ball graph for reliability and area coverage purposes. In particular, we investigate a method for providing more even area coverage without decreasing the reliability of the graph obtained using Algorithm 1 or the simpler version from [8] described in Section 2.1.

### v3

# Vertex Addition to a Ball Graph With Application to Reliability and Area Coverage in Autonomous Swarms

@@ Abstract
Index Terms:
ball graph, disk graph, vertex placement, reliability, area coverage, autonomous swarm, satellite swarm, swarm robotics, formation planning.
I Introduction
@@
One method to keep agents spread out is to ensure some minimum distance between any given pair. The main contribution of this paper is an efficient algorithm to enumerate all possible neighborhoods that a vertex $v$ could have when added to a unit disk graph, under the constraint that no other vertices are within some fixed Euclidean distance $b$ of $v$ (Algorithm 1 of Section III-A). We use this algorithm to move vertices one at a time, under the same constraint, to locations which maximize the reliability of the resulting graph. An alternative method is to use a spring layout- (or Fruchterman-Reingold)-type algorithm, treating vertices like similarly charged particles which repel and edges like springs which attract or repel based on the resting length; we provide a modified spring layout algorithm which avoids destroying edges in a unit ball graph in Section III-B.
The remainder of this paper is organized as follows. First, in Section II, we review an algorithm from our preliminary conference paper [20] to enumerate all possible neighborhoods for a vertex added to a unit ball graph. This can be used to move a vertex in a unit ball graph to a location which maximizes the network reliability of the resulting graph. In Section III, we consider both reliability and area coverage. In Section III-A, we describe our main algorithm in terms of unit disk graphs, noting that this algorithm can be generalized to unit ball graphs as well. In Section III-B, we provide a variation of a classical algorithm of Fruchterman and Reingold [21] to space out vertices in a ball graph without destroying edges. In Section IV, we compare the effectiveness of the algorithms in Sections III-A and III-B in producing unit disk graphs with high reliability and evenly spread vertices over a region. The comparison shows the former algorithm to be particularly effective. We conclude with some open questions and future directions in Section V.
I-A Definitions and notations
@@ I-B Problem statement
II Radial sweeps and ball graphs
We begin with preliminaries on radial sweep algorithms as they apply to disk and ball graphs. We outline an algorithm from our preliminary paper [20], using radial sweeps to list all possible neighborhoods that a new vertex could have when added to a disk or ball graph. This will provide a basis for our main algorithm, described in Section III.
II-A Radial Sweeps
@@
$\theta_{\mathrm{min}}$$\theta_{\mathrm{max}}$vvww
Figure 1: A circle of radius 11 revolves in a counterclockwise motion around a fixed point $v$ on its boundary in a radial sweep. A point $w$ enters and exits the circle at the angles $\theta_{\mathrm{min}}$ and $\theta_{\mathrm{max}}$, respectively. (Adapted from [20].)
Performing these trigonometric calculations for each vertex $v$ in $V$ and each vertex $w$ in VvV-v with $\|v-w\|_{2}<2$, one can determine all subsets $N$ of $V$ for which some radius-$1$ circle contains all of the points in $N$ and none in VNV-N as follows. Each time a point $w\in V$ enters or exits the circle sweeping around a point $v\in V$, we record the subset $S$ of $V$ contained in the interior this circle with $v$ and $w$ on its boundary. By shifting and/or rotating the circle slightly, one can obtain a unit circle with any of $S$, $S\cup v$, $S\cup w$, or $S\cup\{v,w\}$ in its interior. All of these are subsets $N$ of $V$ as described. On the other hand, if $N$ is the subset of $V$ contained in a given radius-$1$ circle, then the circle can be shifted until two points in $V$ lie on its boundary. Thus, after performing a radial sweep around each vertex in $V$, we obtain in this manner all such subsets $N$ of $V$.
@@
II-B Unit disk and unit ball graphs
@@
Given the dynamic nature of swarm formations, it is natural to ask how the communication network changes with the formation. Generally, one might like to know which changes to the graph structure of a unit ball graph $G$ are obtainable by repositioning vertices. This problem is difficult in general, as the recognition problem for unit ball graphs is NP-hard [24]. However, using the algorithm described in Section II-A, we are able to enumerate in polynomial time all of the unit ball graphs obtainable by moving a single vertex in $G$. This can be modeled by deleting a vertex $v$ from $G$ and adding it back in a new location. The algorithm summarized in Section II-A can be used to list all of the possible neighborhoods that $v$ could have.
III Reliability and area coverage
@@
In the context of many a swarm mission, like satellite imaging of a region, such a formation may not be desirable. A natural question then arises: which unit disk graphs have evenly distributed vertices over a given area and have high reliability? In the spirit of our first algorithm, we propose to add or move vertices one at a time, maximizing reliability at each iteration, but imposing the constraint that no vertex be within some fixed distance $b$ of the vertex being added or moved. We present such an algorithm in Section III-A, and we propose an alternative solution to the area coverage problem using a modified spring layout algorithm in Section III-B.
III-A Finding neighborhoods for a vertex with a buffer
@@
There are a number of ways that one might handle this problem. We propose a method whereby a buffer distance is imposed around each vertex as it is added to the unit disk graph $G$. For a value $b$ with $0<b<1$, we note that a vertex which has no neighbors within distance $b$ must have all of its neighbors contained in the $(b,1)$-annulus centered at that vertex (see Figure 2).
@@
Figure 2: The neighbors of a vertex $v$ added with a buffer to a disk graph are contained within an annulus with no vertices in its inner circle. The buffer is depicted by a shaded region, and existing edges in the graph before the addition of $v$ are depicted by dashed edges.
In other words, the possible neighborhoods for a new vertex added with a buffer are precisely the subsets $S$ of the vertex set $V$ of $G$ such that $S$ is contained in a $(b,1)$-annulus and all points in VSV-S lie outside of the radius-$1$ circle. We use a modified radial sweep algorithm to determine all such subsets $S$ of $V$. This subsection is devoted to proving the following.
@@
Let $G$ be a unit disk graph with vertex set $V$, and let $0<b<1$. The list of possible neighborhoods for a vertex $v$ added to $G$ such that $\|v-w\|_{2}>b$ for all $w\in V$ can be obtained in $\mathcal{O}(n\Delta)$ time, where $\Delta$ is the maximum cardinality of a subset of $V$ contained in a radius-$2$ circle.
@@
Algorithm 1 Enumerate neighborhoods for a new vertex added to a unit disk graph with no other vertices within distance $b$ ($0<b<1$).
@@
7:   $L_{1}$ $\leftarrow$ $\textsc{OuterSweep}(V,v,b)$
@@
11:   $L_{2}$ $\leftarrow$ $\textsc{InnerSweep}(V,v,b)$
@@
5:  $L$ $\leftarrow$ $\{\}$
6:  for $w\in V-v$ with $\|v-w\|_{2}<2$ do
7:   $\triangleright$ $w$ enters & exits the sweep
8:   if $1-b<\|v-w\|_{2}<1+b$ then
9:     $\triangleright$ $w$ enters & exits both circles
10:     $\theta_{\mathrm{min}}$ $\leftarrow$ angle $w$ enters outer circle
11:     $\theta_{\mathrm{dis}}$ $\leftarrow$ angle $w$ enters inner circle
12:     $\theta_{\mathrm{re}}$ $\leftarrow$ angle $w$ exits inner circle
13:     $\theta_{\mathrm{max}}$ $\leftarrow$ angle $w$ exits outer circle
14:   else
15:     $\triangleright$ $w$ only enters & exits the outer circle
16:     $\theta_{\mathrm{min}}$ $\leftarrow$ angle $w$ enters outer circle
17:     $\theta_{\mathrm{dis}}$, $\theta_{\mathrm{re}}$ $\leftarrow$ None
18:     $\theta_{\mathrm{max}}$ $\leftarrow$ angle $w$ exits outer circle
19:   end if
20:   add $(w,\theta_{\mathrm{min}},\theta_{\mathrm{dis}},\theta_{\mathrm{re}},\theta_{\mathrm{max}})$ to $L$
21:  end for
22:  return $L$
23:end function
@@
5:  $L$ $\leftarrow$ $\{\}$
6:  for $w\in V-p$ with $\|v-w\|_{2}<1+b$ do
7:   $\triangleright$ $w$ enters & exits the sweep
8:   if $1-b<\|v-w\|_{2}<2b$ then
9:     $\triangleright$ $w$ enters & exits both circles
10:     $\theta_{\mathrm{min}}$ $\leftarrow$ angle $w$ enters outer circle
11:     $\theta_{\mathrm{dis}}$ $\leftarrow$ angle $w$ enters inner circle
12:     $\theta_{\mathrm{re}}$ $\leftarrow$ angle $w$ exits inner circle
13:     $\theta_{\mathrm{max}}$ $\leftarrow$ angle $w$ exits outer circle
14:   else if $\|v-w\|_{2}>2b$ then
15:     $\triangleright$ $w$ only enters & exits outer circle
16:     $\theta_{\mathrm{min}}$ $\leftarrow$ angle $w$ enters outer circle
17:     $\theta_{\mathrm{dis}}$, $\theta_{\mathrm{re}}$ $\leftarrow$ None
18:     $\theta_{\mathrm{max}}$ $\leftarrow$ angle $w$ exits outer circle
19:   else if $\|v-w\|_{2}<1-b$ then
20:     $\triangleright$ $w$ only enters & exits inner circle
21:     $\theta_{\mathrm{min}}$, $\theta_{\mathrm{max}}$ $\leftarrow$ None
22:     $\theta_{\mathrm{dis}}$ $\leftarrow$ angle $w$ enters inner circle
23:     $\theta_{\mathrm{re}}$ $\leftarrow$ angle $w$ exits inner circle
24:   end if
25:   add $(w,\theta_{\mathrm{min}},\theta_{\mathrm{dis}},\theta_{\mathrm{re}},\theta_{\mathrm{max}})$ to $L$
26:  end for
27:  return $L$
28:end function
First, we sweep a $(b,1)$-annulus around a fixed point $v\in V$ on its outer boundary (see Algorithm 2). Consider the $(b,1)$-annulus whose center is at $v+\left[\begin{matrix}1&0\end{matrix}\right]^{\mathsf{T}}$. We imagine a full counterclockwise rotation of the annulus around the fixed point $v$. Points $w\in V$ with $1-b<\|v-w\|_{2}<1+b$ will enter and exit both the inner and outer circles of the annulus as it rotates. Using trigonometric equations similar to those described in Section II-A, we record the angles of rotation for each of these four moments: let $\theta_{\mathrm{min}}$ denote the angle at which $w$ enters the outer circle, $\theta_{\mathrm{dis}}$ the angle at which $w$ enters the inner circle (and disappears from the annulus), $\theta_{\mathrm{re}}$ the angle at which $w$ exits the inner circle (and reappears in the annulus), and $\theta_{\mathrm{max}}$ the angle at which $w$ exits the outer circle. We add the $5$-tuple $(w,\theta_{\mathrm{min}},\theta_{\mathrm{dis}},\theta_{\mathrm{re}},\theta_{\mathrm{max}})$ to a list $L$. Points $w\in V$ with $\|v-w\|_{2}<1-b$ or $\|v-w\|_{2}>1+b$ will enter and exit only the outer circle of the annulus as it rotates about $v$. In this case, we set $\theta_{\mathrm{dis}}=\mathrm{None}$ and $\theta_{\mathrm{re}}=\mathrm{None}$.
@@
5:  $M$ $\leftarrow$ [][\ ]
6:  $R$ $\leftarrow$ $\textsc{InitialSet}(L)$
7:  $\triangleright$ See Algorithm 5 in Appendix B
8:  $B$ $\leftarrow$ $\{\}$
9:  $\triangleright$ Bad vertices: contained in the inner circle
10:  for $(w,\theta_{\mathrm{min}},\theta_{\mathrm{dis}},\theta_{\mathrm{re}},\theta_{\mathrm{max}})\in L$ and  do
11:   for $\theta\in\{\theta_{\mathrm{min}},\theta_{\mathrm{dis}},\theta_{\mathrm{re}},\theta_{\mathrm{max}}\}$ do
12:     if $\theta$ is not None then
13:      add $(w,\theta,\mathrm{type})$ to $M$, where $\mathrm{type}\in\{\min,\mathrm{dis},\mathrm{re},\max\}$
14:     end if
15:   end for
16:   if $\theta_{\mathrm{re}}<\theta_{\mathrm{dis}}$ then
17:     $\triangleright$ $w$ is bad when sweep begins
18:     add $(w,\theta_{\mathrm{dis}},\mathrm{type}=\mathrm{dis})$ to $B$
19:   end if
20:  end for
21:  Sort $M$ by angles $\theta$ in increasing order
22:  $\mathrm{Sets}$ $\leftarrow$ $\{\}$
23:  for $(w,\theta,\mathrm{type})\in M$ do
24:   if $\mathrm{type}=\max$ then remove $w$ from $R$
25:     if $B=\emptyset$ then add $R$ and $R\cup w$ to $\mathrm{Sets}$
@@
27:   else if $\mathrm{type}=\min$ then add $w$ to $R$
28:     if $B=\emptyset$ then add $R$ and RwR-w to $\mathrm{Sets}$
@@
30:   else if $\mathrm{type}=\mathrm{dis}$ then remove $w$ from $R$
31:     if $B=\emptyset$ then add $R\cup w$ to $\mathrm{Sets}$
32:     end if
33:     add $w$ to $B$
34:   else if $\mathrm{type}=\mathrm{re}$ then add $w$ to $R$
35:     remove $w$ from $B$
36:     if $B=\emptyset$ then add $R$ to $\mathrm{Sets}$
37:     end if
38:   end if
39:  end for
40:  return $\mathrm{Sets}$
41:end function
We now describe how to obtain the sets $S$ of vertices contained in a $(b,1)$-annulus with VSV-S outside of its outer boundary (see Algorithm 4). By Proposition 2, it suffices to find the sets $S$ contained in a $(b,1)$-annulus with two points on its boundary (keeping track of whether they are on the outer or inner boundary) and all other points in VSV-S outside of the outer boundary. One of the points on the boundary will be the point $v$ we are sweeping around, and the other will be a point that is entering or exiting the sweep.
@@
From $L_{v}$, we make a new list $M_{v}$ of tuples $(w,\theta,\mathrm{type})$ to record the angles $\theta$, in increasing order, at which points $w$ enter and exit the annulus during the sweep, along with whether it is entering or exiting and which boundary it is passing through. Specifically, $\mathrm{type}$ is one of $\min$’, $\mathrm{dis}$’, $\mathrm{re}$’, or $\max$’ depending on whether $\theta$ is the angle at which $w$ enters the outer circle, enters the inner circle, exits the inner circle, or exits the outer circle of the annulus, respectively.
Now, we go through the tuples in $M_{v}$ in increasing order of $\theta$. If a vertex $w$ exits the outer circle (that is, if the tuple $(w,\theta,\mathrm{type})$ has type $\max$’), we remove $w$ from $R$. We then check if the set of bad vertices $B$ is empty and, if it is, add both $R$ and $R\cup w$ to a list $\mathrm{Sets}$ which will be the output of Algorithm 4. If a vertex $w$ enters the outer circle (if the tuple has type $\min$’), we add it to $R$, and if $B=\emptyset$, we add $R$ and RwR-w to $\mathrm{Sets}$. If a vertex $w$ enters the inner circle (if the tuple has type $\mathrm{dis}$’), we remove $w$ from $R$, check if $B=\emptyset$, and, if it is, add $R\cup w$ to $\mathrm{Sets}$. Then, we add $w$ to $B$. Finally, if a vertex $w$ exits the inner circle (if the tuple has type $\mathrm{re}$’), we add it to $R$ and remove it from $B$. If now $B=\emptyset$, we add $R$ to $\mathrm{Sets}$.
@@
Performing Algorithms 2 and 3 each $n=|V|$ times, by Proposition 2 we obtain all subsets $S$ of $V$ for which there exists a $(b,1)$-annulus containing $S$ and with every point in VSV-S outside of the radius-$1$ circle in $\mathcal{O}(n^{2})$ time. This is summarized in Algorithm 1.
@@
By adjusting the value of $b$, we can use this algorithm to fill in a region with a reliable graph whose vertices are evenly spaced apart. In Section IV, we will compare this algorithm to another method for spreading out vertices without decreasing reliability which we presently describe.
III-B Fruchterman-Reingold for ball graphs
We now describe an alternative approach for modifying a unit ball graph for reliability and area coverage purposes. In particular, we investigate a method for providing more even area coverage without decreasing the reliability of the graph obtained using Algorithm 1 or the simpler version from [20] described in Section II-A.